four empty trucks are on a bridge that has a maximum capacity of 30,000 pounds. If the average (arithmetic mean) weight of three of the trucks is 7,800 pounds, what is the greatest possible safe weight of the fourth truck?
Found 2 solutions by checkley77, nyc_function:
Answer by checkley77(12844)

3*7,800+x=30,000
23,400+x=30,000
x=30,000-23,400
x=6,600 pounds for the 4 th. truck.

Answer by nyc_function(2741)

30,000- 3(7800) = 6600